OpenStudy (aqualungx):

Suppose that a customer service center claims that the telephone call it receives last 112 seconds on average, with a standard deviation of 9 seconds. If you took a sample of 324 telephone calls, which of the following mean times would be within the 95% confidence interval? A. 115.5 seconds B. 112.5 seconds C. 114.5 seconds D. 113.5 seconds

OpenStudy (aqualungx):

@ganeshie8

ganeshie8 (ganeshie8):

margin of error = 1.96*9/sqrt(324) = 0.98 so the confidence interval is\[\large (112-0.98, 112+0.98) = (111.02, 112.98 )\]

ganeshie8 (ganeshie8):

look at the options and pick the one that lies in that interval

OpenStudy (aqualungx):

so it would be 112.5?

ganeshie8 (ganeshie8):

Yep!

OpenStudy (aqualungx):

Thanks a lot!

ganeshie8 (ganeshie8):

yw